OpenStudy (anonymous):

find the indefinite integral of 9/2x+1

OpenStudy (goformit100):



OpenStudy (goformit100):

?

sam (.sam.):

Or is it \[\int\limits \frac{9}{2x+1}dx\]

OpenStudy (anonymous):

YES that is the question.

sam (.sam.):

You can integrate it directly, or by substituting u=2x+1

sam (.sam.):

When \[u=2x+1 \] \[du=2dx\] \[\frac{du}{2}=dx\] \[\int\limits\limits \frac{9}{2x+1}dx\] \[\int\limits\limits \frac{9}{u}\frac{1}{2}du\] \[\frac{9}{2} \int\limits \frac{1}{u}du\] Then finish it up

OpenStudy (anonymous):

Thanks got it, and finished it off . Much appreciated.

sam (.sam.):

Alright gl :)